OpenStudy (anonymous):

what are the real or imaginary solutions of the polynomial equation 125x^3 + 343 = 0

OpenStudy (anonymous):

factor the sum of two cubes.

OpenStudy (watchmath):

\((5x)^3+9^3=0\) \((5x)^3=-9^3\) \(5x=-9\) \(x=-9/5\)

OpenStudy (anonymous):

\[(5x^3)+7^3=(5x+7)((5x)^2-5x\times 7 + 7^2)\]

OpenStudy (watchmath):

oh I forgot the imaginary ...

OpenStudy (anonymous):

and cube root of 343 is 7 not 9 i think

OpenStudy (watchmath):

yes :) you were right sattelite. So ignore my answer :D.

OpenStudy (anonymous):

i am using \[a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (dumbcow):

-7/5 7^3 = 343

OpenStudy (watchmath):

we need two more complex solutions dumbcow

OpenStudy (anonymous):

so we have \[(5x+7)(25x^2-35x+49)=0\] \[5x+7=0\] \[x=-\frac{7}{5}\]

OpenStudy (anonymous):

and the other two you find using the quadratic formula.

OpenStudy (anonymous):

i will write them out if you like

OpenStudy (anonymous):

\[\frac{35\pm \sqrt{35^2-4\times 25\times 49}}{2\times 25}\] \[\frac{35\pm \sqrt{-3675}}{50}\] \[\frac{35\pm 5\sqrt{147}i}{50}\] \[\frac{7\pm \sqrt{147}i}{10}\]